### v1

@@ Abstract
Recent advances in Large Language Models (LLMs) have catalyzed the development of Large Multimodal Models (LMMs). However, existing research primarily focuses on tuning language and image instructions, ignoring the critical pretraining phase where models learn to process textual and visual modalities jointly. In this paper, we propose a new pretraining paradigm for LMMs to enhance the visual comprehension capabilities of LLMs by introducing a novel cross-modal comprehension stage. Specifically, we design a dynamically learnable prompt token pool and employ the Hungarian algorithm to replace part of the original visual tokens with the most relevant prompt tokens. Then, we conceptualize visual tokens as analogous to a “foreign language” for the LLMs and propose a mixed attention mechanism with bidirectional visual attention and unidirectional textual attention to comprehensively enhance the understanding of visual tokens. Meanwhile, we integrate a detailed caption generation task, leveraging rich descriptions to further facilitate LLMs in understanding visual semantic information. After pretraining on 1.5 million publicly accessible data, we present a new foundation model called Croc. Experimental results demonstrate that Croc achieves new state-of-the-art performance on massive vision-language benchmarks. To support reproducibility and facilitate further research, we release the training code and pre-trained model weights at https://github.com/deepglint/Croc.
@@ 1 Introduction
The rapid expansion of mobile networks has accelerated the generation of vast data volumes, presenting unprecedented opportunities for the development and application of Large Language Models (LLMs) (An et al., 2024; Touvron et al., 2023; Bai et al., 2023). Despite their effectiveness, LLMs are primarily confined to processing textual inputs. To expand their multimodal perceptual capabilities, there is an increasing research focus on Large Multimodal Models (LMMs) (Yin et al., 2023; Jin et al., 2024; Yang et al., 2023c), which are designed to process and integrate inputs across multiple modalities.
As a milestone in LMM research, LLaVA (Liu et al., 2024b) leverages the language-only capabilities of GPT-4 (Achiam et al., 2023) to generate multimodal language-image instruction-following datasets, demonstrating impressive multimodal conversational capabilities. Building on this groundwork, LLaVA-1.5 (Liu et al., 2024a) enhances performance through simple modifications to the original LLaVA framework and incorporates academically oriented Visual Question Answering (VQA) datasets with structured response formatting prompts. In parallel, BLIP-2 (Li et al., 2023a) and MiniGPT-4 (Zhu et al., 2023) connect a frozen pre-trained vision encoder and a language model through a trainable Q-Former or a linear layer, effectively mapping image features into the input embedding space of the language model. Nonetheless, these methods achieve only a superficial integration of image features within the language model’s embedding space. In contrast, CogVLM (Wang et al., 2023) introduces a trainable visual expert module into the attention and Feed-Forward Network (FFN) layers of the language model. Despite this innovation, the freezing of the LLM limits its capability to attain an in-depth understanding of visual features directly.
Recent research highlights the crucial role of the pretraining process in LMMs. Flamingo (Alayrac et al., 2022) synergistically integrates pre-trained vision and language models through pretraining on comprehensive multimodal web corpora that combine text and image, enabling the performance of diverse multimodal tasks such as captioning, visual dialogue, and visual question-answering. However, the usage of large-scale pretraining data results in substantial resource consumption. VILA (Lin et al., 2024) reveals that the pretraining process substantively augments various model capabilities, including multi-image reasoning, improved in-context learning, and enriched world knowledge. Nevertheless, VILA uses 50M of the interleaved pre-training corpus to improve data diversity, which is more computationally expensive than LLaVA-1.5. LaVIT (Jin et al., 2023) introduces a meticulously designed visual tokenizer to transform non-linguistic images into a sequence of discrete tokens, thereby rendering them analogous to a foreign language that is interpretable by LLMs. However, this direct input of visual tokens into LLMs and fostering visual understanding through next-token prediction encounters significant hurdles. This limitation primarily originates from the inherent discrepancies between visual and textual tokens, particularly the absence of a robust causal linkage between sequential visual tokens. Furthermore, the application of unidirectional attention mechanisms in this context further restricts the LLM’s capacity to effectively comprehend discrete visual tokens.
@@
In this paper, we introduce a novel pretraining paradigm for LMMs designed to significantly enhance the visual comprehension capabilities of LLMs by incorporating a pioneering cross-modal comprehension stage. Specifically, we design a dynamically learnable prompt token pool and apply the Hungarian algorithm to selectively replace part of the original visual tokens with the most relevant prompt tokens. Then we conceptualize visual tokens as analogous to a “foreign language” for the LLMs and propose a mixed attention mechanism with bidirectional visual attention and unidirectional textual attention to improve the understanding of visual tokens. Meanwhile, we integrate a detailed caption generation task, leveraging rich image descriptions to further facilitate LLMs in understanding visual semantic information. Experiment results demonstrate that our proposed Croc model achieves new state-of-the-art performance across multiple benchmarks. The main contributions of this paper are summarized as follows:
@@
We introduce a new pretraining paradigm to enhance the visual comprehension capabilities of LLMs by introducing a novel cross-modal comprehension stage. This stage integrates visual token reconstruction and targets detailed caption generation.
@@
For visual token reconstruction, we design a dynamically learnable prompt token pool and employ the Hungarian algorithm to replace part of the original image tokens with the most relevant prompt tokens. In addition, we propose a mixed attention mechanism with bidirectional visual attention and unidirectional textual attention for more comprehensive visual token understanding.
@@
Experimental results demonstrate that our proposed Croc model achieves new state-of-the-art across various benchmarks and exhibits robust visual understanding and reasoning capabilities. We will release the training code and models to facilitate future research.
@@ 2 Related Work
Large Multimodal Model Pre-training. As a significant advancement, LLaVA (Liu et al., 2024b) meticulously filters the CC3M (Changpinyo et al., 2021) dataset down to 595K and maintains the frozen state of both the visual encoder and the LLM weights, exclusively training the projection layer to align features from the visual encoder and the LLM. However, this strategy predominantly results in limited deep feature integration between the visual encoder and the LLM, primarily due to the constraints imposed by the projection layer. To address this limitation, CogVLM (Wang et al., 2023) introduces a trainable visual expert module into the attention and feed-forward network layers of the language model. Despite this innovation, constrained by the frozen state of the LLM, it continues to face challenges in comprehending the “foreign language” of visual tokens. Recently, LaVIT (Jin et al., 2023) introduces a well-designed visual tokenizer to convert non-linguistic images into a sequence of discrete tokens. However, directly inputting visual tokens into LLM to enhance visual understanding of LLM through next-token prediction presents significant limitations. VILA (Lin et al., 2024) proposes an interleaved pertraining stage to augment the LLM to support visual input, but it relies on a 50M pertraining dataset, necessitating considerable computational resources.
@@
Cross-Model Comprehension. To improve the performance of Masked Image Modeling (MIM) based pre-training methods, MVP (Wei et al., 2022) initially explores the integration of multimodal pre-training within the MIM framework. Subsequently, diverging from conventional methods that predominantly predict raw pixels or low-level features, MILAN (Hou et al., 2022) adopts an innovative approach by reconstructing image features infused with substantial semantic content derived from caption supervision. UnMasked Teacher (Li et al., 2023b) selectively masks video tokens exhibiting low semantic content and aligns the remaining unmasked tokens through a linear projection to their counterparts from the teacher model. Experimental results confirm that this approach achieves state-of-the-art performance across various video-related tasks. In a recent study, RILS (Yang et al., 2023b) introduces a novel pre-training framework that employs masked visual reconstruction within a language semantic space. This framework facilitates the extraction of structured information by vision models through the accurate semantic prediction of masked tokens. Meanwhile, EVA (Fang et al., 2023) demonstrates that recovering the masked-out tokenized semantic vision features is an efficient strategy for vision-centric representation learning, obviating the need for semantic feature quantization or further tokenization. Inspired by the above works, we propose a visual token reconstruction task to improve the visual comprehension capability of LLMs.
3 Methodology
3.1 Preliminaries of LLaVA and LLaVA-1.5
@@ 3 Methodology
As the seminal work of visual instruction tuning, LLaVA (Liu et al., 2024b) presents the first attempt to use language-only GPT-4 (Achiam et al., 2023) to generate multimodal language-image instruction-following data. The framework of LLaVA comprises three essential components: a Visual Encoder for transforming input images into distinct visual embeddings, a Projector for mapping visual embeddings into the textual embedding space, and a Large Language Model for processing both visual and textual tokens and generating corresponding responses. LLaVA utilizes a two-stage instruction-tuning process. In the first stage, image-text pairs are converted to the single-turn conversation which requests the assistant to describe the image. Given the input visual tokens and textual tokens , both and are fed into LLM to produce a coherent response. The ground-truth prediction answer is represented by the original caption . For a sequence of length , the probability of generating contextually original caption = is calculated as follows:
In the first stage, both the visual encoder and LLM weights are frozen and only the projection layer is updated. In the second stage, LLaVA keeps the visual encoder weights frozen and updates both the pre-trained weights of the projection layer and the LLM. With simple modifications to LLaVA, LLaVA-1.5 (Liu et al., 2024a) integrates CLIP ViT-L/14@336px (Radford et al., 2021) with an MLP projection and incorporates academic task-oriented VQA data with response formatting prompts, resulting in better multimodal comprehension capability.
3.2 Cross-Modal Comprehension
In this section, we propose a novel cross-modal comprehension pre-training method. Fig. 2 illustrates the training pipeline of the proposed Croc model. Unlike LLaVA, Croc includes an additional pre-training phase between stages 1 and 2. We conceptualize image tokens as a foreign language of the LLM and design a dynamically learnable prompt token pool to replace part of the original visual tokens with the most relevant prompt tokens. To facilitate the image token reconstruction, we design a mixed mechanism of bidirectional visual attention and unidirectional textual attention. Meanwhile, we introduce a detailed caption generation task to further enhance the LLM’s understanding of visual tokens.
Prompt Visual Token Generation. Inspired by EVA (Fang et al., 2023), we use LLM to reconstruct the masked visual tokens conditioned on visible image tokens. Given an image , we first extract visual features using the image encoder . Following LLaVA, we select the features before and after the last Transformer layer, and the visual projector translates the visual features into visual tokens . After that, unlike the previous works (He et al., 2022; Li et al., 2023b), we introduce a learnable prompt token pool to replace part of visual tokens with the most relevant prompt tokens. To ensure full utilization of the prompt tokens in the token pool, we use the Hungarian algorithm (Kuhn, 1955) to associate each masked visual token with a corresponding prompt token. We denote the prompt token pool as , where and represent the number of prompt tokens and the feature dimension. Under the mask ratio of , we get the set of masked visual tokens awaiting replacement. We pad with into a set of size . To find a bipartite matching between and , we search for a permutation of elements with the lowest cost:
Mixed Attention Mechanism. Due to the inherent disparities between visual and textual tokens, the causal interactions between visual tokens are significantly weaker than those observed between textual tokens. Inspired by the UNILM (Dong et al., 2019), we introduce a mixed cross-modal attention mechanism. As illustrated in Fig. 3, to enhance pre-training effectiveness and facilitate a more comprehensive assimilation of visual tokens by the LLM, we employ bidirectional attention among visual tokens and unidirectional attention among textual tokens. Such an arrangement not only allows mutual attention among visual tokens but also allows each textual token to attend to all visual tokens. Therefore, this attention configuration significantly enhances the efficacy of LLM in understanding and learning from visual tokens.
@@ 3.2 Cross-Modal Comprehension
Detailed Caption Data. To improve LLM’s understanding of visual tokens, we propose to use detailed caption data for pre-training. As shown in Fig. 4, compared to brief captions, the detailed captions contain a greater wealth of semantic information about image details, thus providing better guidance for LLM to learn visual tokens more effectively.
@@
Training Objectives. To improve the large language model’s ability to learn visual tokens, we introduce two specific objectives: Visual Token Reconstruction (VTR) and Detailed Caption Generation (DCG). In the prompt visual token generation step, we randomly replace some of the original visual tokens with tokens from our prompt token pool, thus obtaining the prompt visual tokens . Then we concatenate the prompt visual tokens with the instruction text tokens and feed them into an LLM to generate a response of length , the probability of generating the contextual response = can be calculated:
@@
After receiving the response , we split the first 576 tokens to compute the visual token reconstruction loss and the remaining tokens to compute the detailed caption generation loss .
The visual token reconstruction loss is calculated as follows:
@@
where represents the index set of replaced visual tokens. Meanwhile, we maximize the likelihood of text tokens by employing the auto-regressive language modeling objective:
The overall training loss is the combination of and :
where is a loss weight used to balance the influence of different losses.

### v2

@@ Abstract
We introduce a novel pretraining stage for Large Multimodal Models (LMMs) aimed at improving their visual comprehension. We approach this problem using the task of generating masked visual tokens. However, we find that this task has some significant limitations in the context of multi-modality learning, which will limit the LLMs efficacy in learning a strong multi-modal representation. To naturally address this problem, we propose to use a dynamically learnable prompt token pool and the Hungarian algorithm for replacing the masked visual tokens with the most relevant visual prompt tokens. Next, we propose a mixed attention mechanism with bidirectional visual attention to encourage more cross-modality collaboration. Finally, we integrate a detailed caption generation task, leveraging rich descriptions to facilitate LLMs in understanding visual semantic information. After pretraining on 1.5 million publicly accessible data, we present a new foundation model called Croc. Croc reaches state-of-the-art results improving over LLaVA-v1.5-7B by , , and on the MMBench, SEED, and LLaVA-Wild benchmarks, respectively. To support reproducibility and facilitate further research, we release the training code and pre-trained model weights at https://github.com/deepglint/Croc.
@@ 1 Introduction
Large Language Models (LLMs) [61, 1] have made significant progress in text understanding and generation, largely due to their foundation in autoregressive causal-Transformer architectures. These models have demonstrated strong scalability with computation and data. However, LLMs are primarily still limited to processing text. To overcome this limitation, there is a growing focus on Large Multimodal Models [66, 26, 65], which aim to process and integrate information across multiple modalities.
A key milestone in LMM research is LLaVA [42], which converts images into token representations using vision encoders such as CLIP [56], and subsequently maps these tokens into LLM space, treating them as language tokens. LLaVA also leverages GPT-4’s language capabilities [1] to generate multimodal language-image instruction-following datasets. Other works [37, 71, 62] have explored more advanced projection architectures, such as QFormers and experts, to improve performance. These approaches highlight the complexity of multimodal tasks, but also the challenges in integrating visual and textual information.
In parallel, there has been increasing emphasis on pretraining strategies for LMMs. Integrating pre-trained vision and language models with multimodal web corpora that combine text and image has proven effective in improving performance across a range of tasks, such as captioning, visual dialogue, and visual question-answering. Notably, the VILA framework [40] demonstrates that pretraining can significantly enhance model capabilities, though it also introduces substantial computational overhead. Another line of research [25] has introduced novel visual tokenizers, which transform non-linguistic images into discrete token sequences, allowing LLMs to interpret them as foreign languages. However, there is still a disparity in the causal interaction between visual and text tokens, which makes cross-modal comprehension still very challenging.
@@
To address these challenges, we propose to introduce a novel pretraining paradigm for LMMs that enhances visual comprehension through a cross-modal reconstruction stage. Our approach involves designing a dynamically learnable prompt token pool, which allows for the selective replacement of certain visual tokens with the most relevant prompt tokens using the Hungarian algorithm. To improve the LLM’s capacity to understand these visual tokens, we then propose a mixed attention mechanism that combines bidirectional visual attention with unidirectional textual attention. Additionally, we integrate a detailed caption generation task, which uses rich image descriptions to facilitate a deeper understanding of visual semantics. Our experimental results show that this method achieves state-of-the-art performance across multiple benchmarks, demonstrating superior visual understanding and reasoning capabilities. In summary, our contributions are the following:
@@
We propose a new pretraining paradigm designed to enhance visual comprehension, incorporating a novel cross-modal comprehension stage that integrates visual token reconstruction and detailed caption generation.
@@
For visual token reconstruction, we design a dynamically learnable prompt token pool and use the Hungarian algorithm for the linear assignment of masked image tokens. We also propose a mixed attention mechanism that combines bidirectional visual attention and unidirectional textual attention for more comprehensive visual token understanding.
@@
We demonstrate that our proposed model achieves new state-of-the-art performance across various benchmarks, outperforming LlaVA-v1.5-7B by up to showing robust visual understanding and reasoning capabilities.
@@ 2 Related Work
Large Multimodal Model Pre-training. LLaVA [42] uses a subset of the CC3M [7] dataset for a more balanced coverage of concepts. Both the visual encoder and LLM are then frozen while the projection layer is trained to align the visual features and language tokens. However, relying solely on this approach will lead to a limited deep feature integration between the visual encoder and the LLM, mainly due to the restrictions imposed by the projection layer. To address this limitation, CogVLM [62] introduce a trainable visual expert module into the attention and feed-forward network layers of the language model. Despite this additional module, the LLM still remains limited by its frozen state and will continue to struggle interpreting the visual tokens. Recently, LaVIT [25] introduced a new visual tokenizer to convert images into a sequence of discrete tokens. However, directly inputting visual tokens into an LLM to enhance visual understanding with next-token prediction still presents significant difficulties. VILA [40] proposes an interleaved pretraining stage to augment the LLM to support visual input, but it relies on a 50M pertraining dataset, requiring considerable computational resources. In contrast, we efficiently enable LLMs to understand visual inputs with just 1.5M pretraining dataset and without any modifications to the underlying LLaVA architecture itself.
@@
Cross-Modal Comprehension. Masked Image Modeling (MIM) [21] is now a common pre-training strategy for improving visual comprehension. Both 4M [54] and MVP [63] propose to integrate this idea in the context of multi-modality learning. In contrast, MILAN [22] propose to reconstruct the image features infused with semantic content derived from caption supervision. UnMasked Teacher [38] selectively masks video tokens exhibiting low semantic content and aligns the remaining unmasked tokens through a linear projection to their counterparts from the teacher model. In a recent study, RILS [64] introduced a novel pre-training framework that employs masked visual reconstruction within a language semantic space. This framework facilitates the extraction of structured information by vision models through the accurate semantic prediction of masked tokens. Meanwhile, EVA [16] demonstrates that reconstructing the masked tokenized semantic vision features is an efficient strategy for vision-centric representation learning, removing the need for semantic feature quantization or any additional tokenization steps. In contrast to all these prior works, we show that masked image modeling has some significant limitations in the context of large multi-modal models. To address these limitations we propose a novel dynamic token pool to enable effective learning at high mask ratios, alongside a detailed semantic caption generation for strong cross-modal collaboration.
3 Cross-Modal Reconstruction
@@
Preliminaries. LLaVA [42] introduces a method for generating multimodal language-image instruction-following data using language-only GPT-4 [1]. Its framework consists of three main components: a Visual Encoder, which converts input images into visual embeddings; a Projector, which maps these visual embeddings into a textual embedding space; and a Large Language Model (LLM) that processes both visual and textual tokens to generate responses. LLaVA uses a two-stage instruction-tuning process. In the first stage, image-text pairs are converted into single-turn conversations where the assistant describes the image. With visual tokens and textual tokens as input, the LLM produces a coherent response, with the original caption serving as the ground-truth answer. Despite the importance of vision-to-text alignment as a pre-training stage, it is insufficient to enable strong visual comprehension [60].
Croc. To address this limitation we propose an additional pre-training stage, which we describe as visual token reconstruction. In Section 3.1, we explain the need for visual token reconstruction to enhance visual comprehension, noting that a simple masked reconstruction of visual tokens is ineffective due to its tendency to produce degenerate or poorly generalizable features. Thus, we propose to extend masked reconstruction in section 3.2 through using a dynamic token pool, that provides the necessary contextual cues for vision token reconstruction. To encourage a better exploration of the token space provided by the token pool, we use Hungarian matching [34]. Finally, we tie all these ideas together in section 3.3 defining the additional pre-training stage and give the full training pipeline in section 3.4.
3.1 Visual Token Reconstruction
The process of visual token reconstruction reflects the way in which LLMs fill in missing words, but here it involves learning to predict visual content based on contextual cues from the visible portion of the image. While the previous image-to-text alignment stage helps the model associate textual representations with visual features, it is insufficient for the model to fully grasp the fine-grained details and spatial relationships of an image. By training on the task of visual token reconstruction, the model can learn to focus on low-level visual features and develop a holistic representation of the image, including texture, shape, and context.
To implement this process, we use the LLM to reconstruct the masked visual tokens conditioned on the visible image tokens and its associated instruction. Given an image , we first extract visual features using the image encoder . Following LLaVA, we then use a visual projector to translate these features into visual tokens . We then mask these tokens before feeding them into the LLM and having a loss for reconstructing the unmasked tokens. This encourages the LLM to learn a robust representation of the full image in order to reconstruct the masked visual tokens correctly. However, while this idea is promising, we find it has a few limitations in practice which we address in the following section.
3.2 Masked Reconstruction Is Not Enough
Masking and reconstructing the visual tokens can lead to poor visual comprehension. If the mask ratio is too high, the model will have insufficient contextual visual information, leading to a poorly robust representation of the image. Conversely, if the mask ratio is too small, the model will learn a degenerate solution using shortcuts such as feature-wise interpolation. We find that this level of fidelity in the masking ratio (see section 4.2) makes its adoption difficult in practice and is in line with observations from prior work [38, 2].
Dynamic Token Pool. Instead of proposing any hand-crafted masking or annealing strategies [2], we propose to rectify the problem using a dynamic token pool. This pool of tokens provides sufficient contextual visual cues at high masking ratios, where each masked token is replaced with a learnable token from the dynamic pool. We then vary the size of this pool to control the level of difficulty for the generation task.
We denote the prompt token pool as , where and represent the number of prompt tokens and the feature dimension respectively. With a masking ratio of , we obtain the set of masked visual tokens . The masked tokens are first padded with into a set of size before awaiting assignment.
Token Assignment. We explore two different options for assigning tokens from the token pool. The first is using a simple nearest neighbors lookup, however, we find that this can lead to an over-dependence on a small subset of the token pool and thus a poor utilization of the entire token space (see section 4.2). Fortunately, this assignment problem has been widely studied in the field of combinatorics, with the Hungarian algorithm [34] providing a polynomial solution with many efficient GPU implementations [55, 14]. Its generality has subsequently seen its adoption in many other practical applications, such as object detection [5].
The objective we have is to find a bipartite matching between and such that the total cost is minimized - where this cost is the L2 distance between the replaced and original visual token. This objective can be solved by finding a permutation of elements that minimizes the following:
As an illustration, the Hungarian algorithm solves this assignment through iteratively subtracting the minimum entries of the rows and columns. We find that this optimisation step is very cheap and accounts for less than 5% of the forward pass wall-clock time during training.
| Method | LLM | Res. | Pretrain | Finetune | VQAv2 | GQA | VizWiz | SciQA-I | TextVQA | Average |
| BLIP-2 | Vicuna-13B | 2242 | 129M | – | 65.0 | 41.0 | 19.6 | 61.0 | 42.5 | 45.8 |
| InstructBLIP | Vicuna-7B | 2242 | 129M | 1.2M | – | 49.2 | 34.5 | 60.5 | 50.1 | – |
| InstructBLIP | Vicuna-13B | 2242 | 129M | 1.2M | – | 49.5 | 33.4 | 63.1 | 50.7 | – |
| IDEFICS-9B | LLaMA-7B | 2242 | 353M | 1M | 50.9 | 38.4 | 35.5 | – | 25.9 | – |
| IDEFICS-80B | LLaMA-65B | 2242 | 353M | 1M | 60.0 | 45.2 | 36.0 | – | 30.9 | – |
| Qwen-VL | Qwen-7B | 4482 | 1.4B† | 50M | 78.8 | 59.3 | 35.2 | 67.1 | 63.8 | 60.8 |
| Qwen-VL-Chat | Qwen-7B | 4482 | 1.4B∗ | 50M | 78.2 | 57.5 | 38.9 | 68.2 | 60.9 | |
| LLaVA-1.5-7B | Vicuna-7B | 3362 | 558K | 665K | 78.5 | 62.0 | 50.0 | 66.8 | 58.2 | 63.1 |
| LLaVA-1.5-13B | Vicuna-13B | 3362 | 558K | 665K | 80.0 | 63.3 | 53.6 | 61.3 | ||
| mPLUG-Owl2 | LLaMA-7B | 2242 | - | - | 79.4 | 56.1 | 68.7 | 58.2 | 63.4 | |
| Croc-7B | Vicuna-7B | 3362 | 558K+1.5M | 665K | 64.2 | 50.0 | 70.1 | 60.4 | 65.0 | |
| Croc-13B | Vicuna-13B | 3362 | 558K+1.5M | 665K | 80.7 | 57.1 | 72.7 | 60.8 | 67.1 | |
3.3 Detailed Caption Generation
@@
Finally, to improve the understanding of visual tokens, we propose to use detailed caption data for pre-training. As shown in Fig. 4, these detailed captions contain a greater wealth of semantic information, thus providing better guidance for jointly learning and utilizing the visual tokens.
We introduce two specific objectives: Visual Token Reconstruction (VTR) and Detailed Caption Generation (DCG).
@@
In the prompt visual token generation step, we randomly remove and then replace some of the original visual tokens with tokens from our prompt token pool, thus obtaining our prompt visual tokens . We then concatenate these prompt visual tokens with the instruction text tokens and feed them into the LLM for generating a response of length .
Improving Cross-Modal Collaboration. Since the causal interactions between visual tokens are significantly weaker than those observed between textual tokens, we rethink the uni-directional attention mechanism during this stage. Inspired by UNILM [15], we propose to use a mixed cross-modal attention mechanism that enables both the vision and text tokens to attend to all the vision tokens (see figure 3). We find that this simple modification can greatly enhances the LLMs visual understanding and can be straightforwardly implemented by just modifying the attention mask.
Training Objectives. After undergoing a forward pass, using this mixed attention strategy, the probability of generating a response = will be given as:
@@
After receiving the response , we reserve the first 576 tokens for computing the visual token reconstruction loss and the remaining tokens to compute the detailed caption generation loss . The visual token reconstruction loss is calculated as follows:
@@
where represents the index set of replaced visual tokens. We maximize the likelihood of text tokens by employing the auto-regressive language modeling objective:
The overall loss is the sum of and :
where is used to balance the influence of different losses.